OpenStudy (anonymous):

Which line is parallel to the line that passes through the points (2, –5) and (–4, 1)?

OpenStudy (texaschic101):

firsy you need to find the slope...do you know how to do that ?

OpenStudy (anonymous):

yes, (y2-y1), (x2-x1)

OpenStudy (texaschic101):

(y2 - y1) / (x2 - x1)....do not forget to divide

OpenStudy (texaschic101):

so find the slope and tell me what you get

OpenStudy (anonymous):

6/-6 ?

OpenStudy (texaschic101):

which reduces to -1 understand ?

OpenStudy (anonymous):

yes, then what's next ?

OpenStudy (texaschic101):

a parallel line will have the same slope. So now we will use this formula : y - y1 = m(x - x1) you can use either of your points and of course m = -1. Does that make sense to you ?

OpenStudy (texaschic101):

let me get you started... y - y1 = m(x - x1) using slope(m) = -1 and points (2,-5) y - (-5) = -1(x - 2) do you see how I got that ?

OpenStudy (anonymous):

yes

OpenStudy (texaschic101):

lets continue... y - (-5) = -1(x - 2) y + 5 = -x + 2 we now want to get y by itself, so we subtract 5 from both sides y = -x + 2 - 5 combine like terms y = -x - 3 understand ?

OpenStudy (anonymous):

yes, what's next

OpenStudy (texaschic101):

that is your equation, unless it needs to be in Ax + By = C form.....does it ?

OpenStudy (anonymous):

i don't think so because it's just asking what y equals, but y=-x-3 is not an option

OpenStudy (texaschic101):

let me try your other points...hold on

OpenStudy (anonymous):

the options are; y= -x-5 y= -2/3x+3 y= 2/3x-2 y= x+5

OpenStudy (texaschic101):

y - y1 = m(x - x1) using slope -1 and points (-4,1) y - 1 = -1(x -(-4) y - 1 = -1(x + 4) y - 1 = -x - 4 y = -x - 4 + 1 y = -x - 3.....same answer let me check the slope m = (y2 - y1) / (x2 - x1) m = 1 - (-5) / (-4 - 2) m = (1 + 5) / -6 m = 6/-6 = -1 hmmm....are you sure those are the right answer choices ?

OpenStudy (texaschic101):

Is the question written correctly ? It is asking for parallel line and not perpendicular, right ?

OpenStudy (anonymous):

"Which line is parallel to the line that passes through the points " is what the question is asking exactly. That's why I'm confused too

OpenStudy (texaschic101):

The points are written correctly ?

OpenStudy (anonymous):

Here is the question exactly: Which line is parallel to the line that passes through the points (2, –5) and (–4, 1)? A. y= -x-5 B. y= -2/3x+3 C. y= 2/3x-2 D. y= x+5

OpenStudy (texaschic101):

I do not see anything wrong with the way I did it. @gypsy1274 ......please check this

OpenStudy (anonymous):

yes, please. I really need help with this question

OpenStudy (texaschic101):

dont worry...gypsy is good

OpenStudy (anonymous):

I hope so. Thank you so much by the way!!!

OpenStudy (anonymous):

The only thing you need to find here is the parallel line so the only value you should be concerned with is the slope.

OpenStudy (texaschic101):

the slope is -1, but jaytag needs the line parallel .......is the -3 wrong

OpenStudy (anonymous):

The slope is -1 so you are looking for a line with a slope of -1. In the form y=mx+b -> m=slope

OpenStudy (texaschic101):

I see.....jaytag, which answer has a -1 as the slope ?

OpenStudy (anonymous):

You are doing too much work. The y-intercept is irrelevant in this question.

OpenStudy (texaschic101):

thank you so much Gypsy, I knew you could straighten this out

OpenStudy (anonymous):

So, the answer would be A. y=-x-5?

OpenStudy (texaschic101):

yes....thanks to Gypsy

OpenStudy (anonymous):

Thank you both so much!!!

OpenStudy (texaschic101):

sorry...I probably got you really confused

OpenStudy (anonymous):

@texaschic101 There was nothing wrong with what you did, you just went a few more steps than necessary. :-)

OpenStudy (texaschic101):

I have a problem with making things harder then they really are :(